A CYCLIST RIDING WITH ASPEED OF 27 KMPH . AS HE APPROACHES A CIRCULAR TURN ON THE ROAD HE APPLIES BRAKES AND REDUCES HIS SPEED AT CONSTANT RATE OF 0.5 M/S^2,IF RAIUS OF CIRCLE IS 80 M. WHAT IS THE MAGNITUDE AND DIRECTION OF NET ACCELARATIOON.

Grade:11

2 Answers

Akash Kumar Dutta
10 years ago

magnitude of tangential acc=.5 m/s^2
magnitude of centripetal acc.= v^2/r=45/64 m/s^2
then find mag. of net acc as root( tang. acc^2 + centripetal acc. ^2 )
and direction as tan inverse ( 32/45) with the radius of the circle.

KS
10 years ago

You must consider both , Centripetal and tangential Accelerations and for the final answer take the resultant of them.